A certain gas effuses 4 times as fast as oxygen gas. What is the molar mass of the unknown gas
1
Expert's answer
2019-08-19T02:05:58-0400

The equation for the effusion rate is:

"\frac{ER_{1}}{ER_{2}}=(\frac{M_{2}}{M_{1}})^{1/2}"

,where ER is the effusion rate.

Here we have "\frac{ER(X)}{ER(O_{2})}=4" and the molar mass of the oxygen gas is equal to 32 g/mol.

Thus, "\frac{ER(X)}{ER(O_{2})}=(\frac{M(O_{2})}{M(X)})^{1/2}"

Using this formula, taking its square and slightly rearrange it, we could get the following expression "M(X)=M(O_{2})*(\frac{ER(O_{2})}{ER(X)})^2"

and finally "M(X)=32g/mol*(\frac{1}{4})^2=2g/mol"

It is obviously, that the unknown gas is hydrogen gas or H2.